OpenStudy (anonymous):

The product of two consecutive positive integers is 1 more than their sum. Find the integers

OpenStudy (lgbasallote):

let x be first number x+1 = consecutive number x(x+1) = 1 + (x + x + 1) x^2 + x = 2 + 2x x^2 - x -2 = 0 (x-2)(x+1) = 0 x = 2 x = -1 but since you're looking for positive..we take x = 2 so x+ 1 = 3

OpenStudy (anonymous):

Thanks